# The Race between Technology and Woman:

### v8

@@ 6.1.2 Robustness checks
Table 2 presents the robustness of the estimates to controlling for the influence of labor market institutions, product market imperfections, and disembodied factor-biased technological change. The first three rows report the estimates of the substitution parameters after controlling for the influence of labor market institutions. We allow for the possibility that the actual wage may deviate from the competitive wage depending on the collective bargaining coverage, the strictness of employment protection legislation, and the presence and level of minimum wages. Specifically, we add the collective bargaining coverage and the employment protection legislation in log-difference form and the presence of minimum wages and its interaction with their level in first-difference form to equations (32)–(36). The relative magnitude of the estimated substitution parameters remains unchanged with the additional controls. Although the first three rows of Table 2 report the estimates when the three sets of variables are added individually to avoid the multicollinearity problem, the relative magnitude of the estimated substitution parameters remains unchanged even if they are added jointly.
@@
Collective bargaining –0.626 0.310 0.621 0.713 0.681 0.681 0.681 0.681
   coverage (0.449) (0.112) (0.087) (0.109) (0.080) (0.080) (0.080) (0.080)
Employment protection –0.809 0.188 0.568 0.728 0.509 0.509 0.509 0.509
   legislation (0.350) (0.054) (0.070) (0.090) (0.083) (0.083) (0.083) (0.083)
Minimum wages –0.719 0.407 0.615 0.766 0.597 0.597 0.597 0.597
(0.499) (0.082) (0.101) (0.149) (0.078) (0.078) (0.078) (0.078)
Product market –0.899 0.208 0.569 0.748 0.525 0.525 0.525 0.525
   imperfection (0.380) (0.065) (0.048) (0.059) (0.092) (0.092) (0.092) (0.092)
Disembodied factor-biased –0.740 0.322 0.537 0.844 0.677 0.677 0.677 0.677
   technological change (0.171) (0.060) (0.029) (0.035) (0.053) (0.053) (0.053) (0.053)
@@
Collective bargaining –0.726 0.282 0.411 0.599 –0.489 0.265 0.580 0.816
   coverage (0.356) (0.158) (0.106) (0.089) (0.196) (0.089) (0.044) (0.069)
Employment protection –1.410 0.141 0.327 0.584 –0.545 0.262 0.573 0.840
   legislation (0.406) (0.112) (0.068) (0.060) (0.168) (0.066) (0.046) (0.065)
Minimum wages –1.067 0.252 0.389 0.643 –0.431 0.358 0.603 0.880
(0.278) (0.096) (0.076) (0.081) (0.146) (0.087) (0.043) (0.063)
Product market –1.176 0.105 0.344 0.584 –0.618 0.219 0.521 0.803
   imperfection (0.302) (0.073) (0.064) (0.059) (0.168) (0.074) (0.033) (0.046)
Disembodied factor-biased –0.655 0.431 0.700 0.903 –0.352 0.479 0.630 0.903
   technological change (0.230) (0.123) (0.078) (0.048) (0.109) (0.086) (0.049) (0.044)
@@
The fourth row of Table 2 reports the estimates of the substitution parameters after taking into account imperfect competition in the product market. We can relax the assumption of competitive markets by recalculating the rental price of capital based on the external rate of return because all estimating equations hold irrespective of the degree of markup. The estimates of the substitution parameters remain essentially unchanged.
@@ 6.2.1 Elasticities of substitution
Table 3 presents the estimates of the elasticities of substitution in the aggregate production and cost functions. The estimated elasticities of substitution between ICT capital and the four types of labor are greater in the order of male unskilled, female unskilled, male skilled, and female skilled labor in both cases. The relative magnitude of the estimated elasticities are consistent with the results in Table 4.3. The elasticity estimates imply that the expansion of ICT equipment would narrow the gender wage gap and widen the skill wage gap. In this sense, the estimates are consistent with the literature that points to technological advances as causes of changes in the structure of wages and employment. The magnitude of the factor complementarity effects associated with a rise in ICT capital is directly proportional to the differences in substitution elasticities as well as the rate of increase in ICT capital. The elasticity estimates imply that the effects of ICT capital on the gender (skill) wage gap would be greater for skilled (female) labor than for unskilled (male) labor. All else held constant, a one percent increase in ICT capital would reduce the skilled (unskilled) gender wage gap by 0.30 (0.02) percent and raise the male (female) skill wage gap by 0.08 (0.36) percent. At the same time, the magnitude of the relative quantity effects is inversely proportional to substitution elasticities and directly proportional to the rates of increases in the quantities of labor. The elasticity estimates imply that the effects of labor quantities would be greater in the order of female skilled, male skilled, female unskilled, and male unskilled labor.
@@
$k_{i}$ 0. 77 1. 53 2. 50 5. 91 0. 99 $k_{i}$ 0. 79 1. 53 2. 51 6. 21 0. 99
(0. 04) (0. 09) (0. 15) (0. 96) (0. 00) (0. 04) (0. 09) (0. 15) (1. 23) (0. 00)
$\ell_{fh}$ 0. 76 1. 54 2. 48 6. 63 0. 99 $\ell_{fh}$ 0. 75 1. 52 2. 48 6. 29 0. 99
(0. 04) (0. 09) (0. 15) (1. 97) (0. 00) (0. 04) (0. 09) (0. 14) (1. 29) (0. 00)
$\ell_{mh}$ 0. 99 1. 06 2. 51 6. 14 0. 99 $\ell_{mh}$ 1. 12 1. 20 2. 52 6. 21 0. 99
(0. 05) (0. 05) (0. 15) (1. 15) (0. 00) (0. 05) (0. 07) (0. 15) (1. 23) (0. 00)
$\ell_{fu}$ 1. 05 1. 12 1. 96 6. 24 0. 99 $\ell_{fu}$ 1. 31 1. 41 2. 07 6. 25 0. 99
(0. 05) (0. 06) (0. 10) (1. 26) (0. 00) (0. 06) (0. 08) (0. 11) (1. 25) (0. 00)
$\ell_{mu}$ 1. 08 1. 17 2. 28 3. 50 0. 99 $\ell_{mu}$ 1. 71 1. 88 3. 18 4. 29 0. 98
(0. 05) (0. 06) (0. 14) (0. 30) (0. 00) (0. 16) (0. 21) (0. 41) (0. 65) (0. 00)
$k_{o}$ 1. 03 1. 10 1. 76 2. 03 1. 73 $k_{o}$ 1. 42 1. 55 2. 39 3. 02 3. 66
(0. 04) (0. 05) (0. 07) (0. 07) (0. 04) (0. 10) (0. 13) (0. 23) (0. 35) (0. 64)
Notes: The elasticities of substitution are computed using equations (13) and (14) and evaluated at the sample means across all countries and years. Standard errors in parentheses are computed using bootstrap with 500 replications.
The estimated elasticities of labor–labor substitution in the aggregate production (cost) function range from 1.1 to 6.6 (1.2 to 6.3). The elasticity estimates are within the range of standard estimates in the literature, although they are not strictly comparable to those in other studies because of differences in the degree of classification of labor. The estimated aggregate elasticities of substitution between female skilled labor and the other three types of labor are greater in the order of male unskilled, female unskilled, and male skilled labor. Both female skilled and unskilled labor are more complementary to male skilled labor than to male unskilled labor.
The Morishima (1967) elasticities of substitution are asymmetric by definition. The differences in the estimated elasticities of relative demand for capital with respect to wages are greater than those in the estimated elasticities of relative demand for labor with respect to the rental prices of capital. However, in both cases, ICT capital is more complementary not only to skilled labor than unskilled labor but also to female labor than male labor.
The estimated elasticity of substitution between ICT capital and male unskilled labor in the aggregate cost (production) function exceeds one at $\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}=1.71$ ($\epsilon_{\ell_{mu}k_{i}}^{\mathcal{F}}=1.08$). ICT capital is consistently more substitutable with male unskilled labor than with the other three types of labor. The estimates suggest that the income share of male unskilled labor is likely to decline with a fall in the price or a rise in the quantity of ICT capital.
@@ 6.2.2 Contributions of factor inputs
Table 4 presents the quantitative contributions of specific factor inputs to gender and skill premia. The first two columns report the actual and predicted changes in gender and skill premia. The rate of decline in the gender wage gap is greater for unskilled workers than for skilled workers, while the rate of increase in the skill wage gap is greater for male workers than for female workers. From this point of view, the observed pattern of changes in gender and skill premia is consistent with the pattern predicted from the model.
@@
Data Model FCE RQE $\Delta\ln k_{i}$ $\Delta\ln\ell_{fh}$ $\Delta\ln\ell_{mh}$ $\Delta\ln\ell_{fu}$ $\Delta\ln\ell_{mu}$ $\Delta\ln k_{o}$
$\Delta\ln\left(\frac{w_{mh}}{w_{fh}}\right)$ 0. 012 0. 066 –0. 889 0. 955 –0. 888 1. 423 –0. 468 –0. 001 0. 000 0. 000
(0. 050) (0. 069) (0. 049) (0. 069) (0. 071) (0. 028) (0. 001) (0. 000) (0. 000)
$\Delta\ln\left(\frac{w_{mu}}{w_{fu}}\right)$ –0. 057 –0. 110 –0. 170 0. 059 –0. 072 –0. 047 –0. 051 0. 058 0. 001 0. 001
(0. 016) (0. 014) (0. 005) (0. 005) (0. 005) (0. 004) (0. 005) (0. 000) (0. 000)
$\Delta\ln\left(\frac{w_{mh}}{w_{mu}}\right)$ 0. 118 0. 059 0. 377 –0. 318 0. 224 0. 130 –0. 317 0. 023 –0. 001 0. 000
(0. 020) (0. 023) (0. 020) (0. 014) (0. 010) (0. 019) (0. 002) (0. 000) (0. 000)
$\Delta\ln\left(\frac{w_{fh}}{w_{fu}}\right)$ 0. 049 –0. 118 1. 140 –1. 258 1. 039 –1. 340 0. 100 0. 082 0. 000 0. 000
(0. 046) (0. 071) (0. 062) (0. 068) (0. 065) (0. 013) (0. 005) (0. 000) (0. 000)
@@
Notes: The first and second columns report the actual and predicted changes in the logarithms of relative wages. The predicted change is the sum of the factor complementarity effect (FCE) in the third column and the relative quantity effect (RQE) in the fourth column, or equivalently, the sum of the contributions of each factor in the fifth to tenth columns. The figures in the third to tenth columns are computed using equation (19). Standard errors in parentheses are computed using bootstrap with 500 replications.
@@
The third and fourth columns report the changes attributable to the factor complementarity effect and the relative quantity effect. The fifth to tenth columns report the changes attributable to each factor input. This period shows a substantial increase in ICT capital and female skilled labor, a moderate increase in non-ICT capital and male skilled labor, a small increase in female unskilled labor, and a small decrease in male unskilled labor. The increase in ICT capital contributes to narrowing the gender wage gap and widening the skill wage gap. Meanwhile, the increase in male and female skilled labor contributes to widening the gender wage gap and narrowing the skill wage gap. Consequently, whether the gender (skill) wage gap would narrow (widen) depends on the outcome of the race between progress in ICT and advances in female employment (educational attainment).
[…]
@@
Changes in the labor share can be decomposed into those in the income shares of the four types of labor $\ell_{f}$ for $f\in\mathscr{F}=\left\{mh,fh,mu,fu\right\}$, as shown in equation (18).
$\Delta\ln\left(\frac{\sum_{f\in\mathscr{F}}\sum_{n}w_{fn}\ell_{fn}}{py}\right)=\sum_{f\in\mathscr{F}}\left(\frac{\sum_{n}w_{fn}\ell_{fn}}{\sum_{f\in\mathscr{F}}\sum_{n}w_{fn}\ell_{fn}}\right)\Delta\ln\left(\frac{\sum_{n}w_{fn}\ell_{fn}}{py}\right),$ (38)
where $py=\sum_{f\in\mathscr{F}}\sum_{n}w_{fn}\ell_{fn}+\sum_{n}r_{in}k_{in}+\sum_{n}r_{on}k_{on}$. We evaluate the weights on the rates of changes in the income shares of each type of labor ($\sum_{n}w_{fn}\ell_{fn}/\sum_{f\in\mathscr{F}}\sum_{n}w_{fn}\ell_{fn}$) at the sample means across all countries and years. Changes in the labor share between the years 1980 and 2005 can be decomposed as
[…]
$=$ $0.075+0.081-0.200-0.031$
@@
Data Model FCE RQE $\Delta\ln k_{i}$ $\Delta\ln\ell_{fh}$ $\Delta\ln\ell_{mh}$ $\Delta\ln\ell_{fu}$ $\Delta\ln\ell_{mu}$ $\Delta\ln k_{o}$
$\Delta\ln\left(\frac{w_{mh}\ell_{mh}}{py}\right)$ 0. 457 0. 442 0. 148 0. 294 0. 092 0. 050 0. 313 –0. 019 0. 003 0. 003
(0. 016) (0. 021) (0. 011) (0. 016) (0. 011) (0. 017) (0. 002) (0. 000) (0. 000)
$\Delta\ln\left(\frac{w_{fh}\ell_{fh}}{py}\right)$ 1. 230 1. 161 1. 036 0. 125 0. 980 0. 134 0. 059 –0. 018 0. 003 0. 002
(0. 047) (0. 066) (0. 054) (0. 070) (0. 063) (0. 013) (0. 002) (0. 000) (0. 000)
$\Delta\ln\left(\frac{w_{mu}\ell_{mu}}{py}\right)$ –0. 389 –0. 346 –0. 263 –0. 083 –0. 131 –0. 080 –0. 092 –0. 042 –0. 003 0. 003
(0. 011) (0. 008) (0. 010) (0. 004) (0. 004) (0. 003) (0. 001) (0. 000) (0. 000)
$\Delta\ln\left(\frac{w_{fu}\ell_{fu}}{py}\right)$ –0. 122 –0. 024 –0. 084 0. 060 –0. 059 –0. 033 –0. 041 0. 104 0. 003 0. 002
(0. 010) (0. 010) (0. 007) (0. 005) (0. 003) (0. 003) (0. 004) (0. 000) (0. 000)
@@
Notes: The first and second columns report the actual and predicted changes in the logarithms of labor shares. The predicted change is the sum of the factor complementarity effect (FCE) in the third column and the relative quantity effect (RQE) in the fourth column, or equivalently, the sum of the contributions of each factor in the fifth to tenth columns. The figures in the third to tenth columns are computed using equation (16). Standard errors in parentheses are computed using bootstrap with 500 replications.
@@
Table 5 presents the quantitative contributions of specific factor inputs to the income shares of the four types of labor. The first two columns report the actual and predicted changes in the income shares of the four types of labor. The actual changes in the income shares of the four types of labor align with the changes predicted from the model. The predicted changes in the income shares of male and female skilled labor are positive and statistically significant, while those in the income shares of male and female unskilled labor are negative and statistically significant.
@@
The third and fourth columns report the changes attributable to the factor complementarity effect and the relative quantity effect. The two effects work in the same direction for the income shares of male and female skilled labor and male unskilled labor. The former is much greater than the latter for the income shares of female skilled labor and male unskilled labor, although the converse is true for the income share of male skilled labor. The two effects work in opposite directions for the income share of female unskilled labor. However, the direction of the factor complementarity effect aligns with that of the actual change for the income shares of all types of labor. Overall, the factor complementarity effect is essential for understanding changes in the income shares of different types of labor.

### v9

@@ 6.1.2 Robustness checks
Table 3 presents the robustness of the estimates to controlling for the influence of labor market institutions, goods market imperfections, and disembodied factor-biased technological change. The first three rows report the estimates of the substitution parameters after controlling for the influence of labor market institutions. We allow for the possibility that the actual wage may deviate from the competitive wage depending on the collective bargaining coverage, the strictness of employment protection legislation, and the presence and level of minimum wages. Specifically, we add the collective bargaining coverage and the employment protection legislation in log-difference form and the presence of minimum wages and its interaction with their level in first-difference form to equations (28)–(32). The relative magnitude of the estimated substitution parameters remains unchanged with the additional controls. Although the first three rows of Table 3 report the estimates when the three sets of variables are added individually to avoid the multicollinearity problem, the relative magnitude of the estimated substitution parameters remains unchanged even if they are added jointly.
The fourth row of Table 3 reports the estimates of the substitution parameters after taking into account imperfect competition in goods markets. We can relax the assumption of competitive markets by recalculating the rental price of capital based on the external rate of return because all estimating equations hold irrespective of the degree of markup. The estimates of the substitution parameters remain essentially unchanged.
@@
Collective bargaining –0. 601 0. 307 0. 626 0. 720 0. 676 0. 676 0. 676 0. 676
   coverage (0. 408) (0. 074) (0. 047) (0. 065) (0. 074) (0. 074) (0. 074) (0. 074)
Employment protection –0. 855 0. 166 0. 553 0. 707 0. 496 0. 496 0. 496 0. 496
   legislation (0. 328) (0. 056) (0. 055) (0. 065) (0. 082) (0. 082) (0. 082) (0. 082)
Minimum wages –0. 668 0. 387 0. 635 0. 784 0. 592 0. 592 0. 592 0. 592
(0. 439) (0. 057) (0. 053) (0. 094) (0. 071) (0. 071) (0. 071) (0. 071)
Goods market –0. 928 0. 191 0. 559 0. 733 0. 518 0. 518 0. 518 0. 518
   imperfections (0. 368) (0. 061) (0. 049) (0. 058) (0. 087) (0. 087) (0. 087) (0. 087)
Disembodied factor-biased –0. 722 0. 336 0. 542 0. 854 0. 674 0. 674 0. 674 0. 674
   technological change (0. 141) (0. 047) (0. 028) (0. 030) (0. 060) (0. 060) (0. 060) (0. 060)
@@
Collective bargaining –0. 891 0. 204 0. 375 0. 566 –0. 458 0. 282 0. 584 0. 809
   coverage (0. 421) (0. 139) (0. 083) (0. 054) (0. 186) (0. 074) (0. 040) (0. 064)
Employment protection –1. 344 0. 152 0. 348 0. 594 –0. 563 0. 246 0. 567 0. 828
   legislation (0. 400) (0. 113) (0. 054) (0. 042) (0. 159) (0. 065) (0. 041) (0. 060)
Minimum wages –1. 117 0. 211 0. 388 0. 636 –0. 418 0. 361 0. 604 0. 872
(0. 365) (0. 109) (0. 066) (0. 063) (0. 139) (0. 081) (0. 040) (0. 059)
Goods market –1. 200 0. 076 0. 330 0. 565 –0. 626 0. 212 0. 518 0. 791
   imperfections (0. 363) (0. 088) (0. 064) (0. 054) (0. 169) (0. 073) (0. 034) (0. 047)
Disembodied factor-biased –0. 661 0. 439 0. 705 0. 908 –0. 371 0. 473 0. 627 0. 899
   technological change (0. 246) (0. 132) (0. 081) (0. 050) (0. 095) (0. 086) (0. 049) (0. 043)
@@ 6.2.1 Elasticities of substitution
Table 4 presents the estimates of the elasticities of substitution in the aggregate production and cost functions. The estimated elasticities of substitution between ICCT capital and the four types of labor are greater in the order of male unskilled, female unskilled, male skilled, and female skilled labor in both cases. The elasticity estimates imply that the expansion of ICCT equipment would narrow the gender wage gap and widen the skill wage gap. In this sense, the estimates are consistent with the literature that points to technological advances as causes of changes in the structure of wages and employment (Autor, Levy and Murnane, 2003; Black and Spitz-Oener, 2010). The magnitude of the factor complementarity effects associated with a rise in ICCT capital is directly proportional to the differences in the elasticities of substitution between ICCT capital and the four types of labor as well as the rate of increase in ICCT capital. The elasticity estimates imply that the effects of ICCT capital on the gender (skill) wage gap would be greater for skilled (female) labor than for unskilled (male) labor. All else held constant, a one percent increase in ICCT capital would reduce the skilled (unskilled) gender wage gap by 0.30 (0.02) percent and raise the male (female) skill wage gap by 0.08 (0.35) percent. At the same time, the magnitude of the relative quantity effects is inversely proportional to the elasticities of labor–labor substitution and directly proportional to the rates of increases in the quantities of labor. The elasticity estimates imply that the effects of labor quantities would be greater in the order of female skilled, male skilled, female unskilled, and male unskilled labor.
@@
$k_{i}$ 0. 78 1. 52 2. 49 5. 66 1. 00 $k_{i}$ 0. 79 1. 53 2. 50 5. 86 1. 00
(0. 04) (0. 08) (0. 14) (0. 82) (0. 00) (0. 04) (0. 09) (0. 14) (0. 99) (0. 00)
$\ell_{fh}$ 0. 76 1. 54 2. 48 6. 22 1. 00 $\ell_{fh}$ 0. 75 1. 52 2. 47 5. 93 1. 00
(0. 04) (0. 09) (0. 14) (1. 44) (0. 00) (0. 04) (0. 08) (0. 13) (1. 04) (0. 00)
$\ell_{mh}$ 0. 99 1. 06 2. 50 5. 80 1. 00 $\ell_{mh}$ 1. 12 1. 20 2. 51 5. 86 1. 00
(0. 04) (0. 05) (0. 14) (0. 93) (0. 00) (0. 05) (0. 07) (0. 14) (0. 99) (0. 00)
$\ell_{fu}$ 1. 05 1. 12 1. 96 5. 90 1. 00 $\ell_{fu}$ 1. 31 1. 41 2. 07 5. 89 1. 00
(0. 05) (0. 05) (0. 10) (1. 01) (0. 00) (0. 06) (0. 07) (0. 10) (1. 01) (0. 00)
$\ell_{mu}$ 1. 07 1. 16 2. 26 3. 44 1. 00 $\ell_{mu}$ 1. 67 1. 83 3. 07 4. 11 1. 00
(0. 05) (0. 06) (0. 13) (0. 27) (0. 00) (0. 14) (0. 18) (0. 34) (0. 53) (0. 00)
$k_{o}$ 1. 03 1. 10 1. 76 2. 03 1. 73 $k_{o}$ 1. 40 1. 53 2. 33 2. 93 3. 49
(0. 04) (0. 05) (0. 07) (0. 07) (0. 04) (0. 09) (0. 11) (0. 19) (0. 29) (0. 51)
Notes: The elasticities of substitution are computed using equations (12) and (13) of Lemma 1 and evaluated at the sample means across all countries and years. Standard errors in parentheses are computed using bootstrap with 500 replications.
The estimated elasticities of labor–labor substitution in the aggregate production (cost) function range from 1.1 to 6.2 (1.2 to 5.9). The elasticity estimates are within the range of standard estimates in the literature, although they are not strictly comparable to those in other studies, such as Acemoglu, Autor and Lyle (2004) and Johnson and Keane (2013), because of differences in the degree of classification of labor. The estimated aggregate elasticities of substitution between female skilled labor and the other three types of labor are greater in the order of male unskilled, female unskilled, and male skilled labor. Both female skilled and unskilled labor are more complementary to male skilled labor than to male unskilled labor.
The Morishima (1967) elasticities of substitution are asymmetric by definition. The differences in the estimated elasticities of relative demand for capital with respect to wages are greater than those in the estimated elasticities of relative demand for labor with respect to the rental prices of capital. However, in both cases, ICCT capital is more complementary not only to skilled labor than unskilled labor but also to female labor than male labor.
The estimated elasticity of substitution between ICCT capital and male unskilled labor in the aggregate cost (production) function exceeds one at $\epsilon_{\ell_{mu}k_{i}}^{\mathcal{C}}=1.67$ ($\epsilon_{\ell_{mu}k_{i}}^{\mathcal{F}}=1.07$). ICCT capital is consistently more substitutable with male unskilled labor than with the other three types of labor. The estimates suggest that the income share of male unskilled labor is likely to decline with a fall in the price or a rise in the quantity of ICCT capital.
@@ 6.2.2 Contributions of factor inputs
Table 5 presents the quantitative contributions of specific factor inputs to changes in the gender premium by skill and the skill premium by gender. The first column reports the observed changes in gender and skill premia, while the second column reports the sum of the changes attributable to factor inputs. These results are obtained without adjusting any parameters to match relative wages in the aggregate economy. The rate of decline in the gender wage gap is greater among unskilled workers than among skilled workers, while the rate of increase in the skill wage gap is greater among male workers than among female workers. From this point of view, the observed pattern of changes in gender and skill premia is consistent with the total contribution of factor inputs.
@@
Data Total FCE RQE $\Delta\ln k_{i}$ $\Delta\ln\ell_{fh}$ $\Delta\ln\ell_{mh}$ $\Delta\ln\ell_{fu}$ $\Delta\ln\ell_{mu}$ $\Delta\ln k_{o}$
$\Delta\ln\left(\frac{w_{mh}}{w_{fh}}\right)$ 0. 012 0. 070 –0. 885 0. 954 –0. 884 1. 423 –0. 469 –0. 001 0. 000 0. 000
(0. 050) (0. 070) (0. 047) (0. 070) (0. 067) (0. 027) (0. 001) (0. 000) (0. 000)
$\Delta\ln\left(\frac{w_{mu}}{w_{fu}}\right)$ –0. 057 –0. 103 –0. 164 0. 061 –0. 070 –0. 045 –0. 049 0. 059 0. 001 0. 000
(0. 015) (0. 013) (0. 005) (0. 005) (0. 004) (0. 004) (0. 004) (0. 000) (0. 000)
$\Delta\ln\left(\frac{w_{mh}}{w_{mu}}\right)$ 0. 118 0. 052 0. 373 –0. 321 0. 222 0. 129 –0. 320 0. 022 –0. 001 0. 000
(0. 019) (0. 023) (0. 019) (0. 014) (0. 010) (0. 019) (0. 002) (0. 000) (0. 000)
$\Delta\ln\left(\frac{w_{fh}}{w_{fu}}\right)$ 0. 049 –0. 121 1. 136 –1. 257 1. 036 –1. 340 0. 100 0. 082 0. 000 0. 000
(0. 046) (0. 071) (0. 059) (0. 069) (0. 062) (0. 013) (0. 005) (0. 000) (0. 000)
$\Delta\ln\left(\frac{\overline{w}_{M}}{\overline{w}_{F}}\right)$ –0. 083 –0. 111 –0. 307 0. 196 –0. 228 0. 147 –0. 094 0. 063 0. 001 0. 000
(0. 020) (0. 022) (0. 012) (0. 017) (0. 014) (0. 006) (0. 004) (0. 000) (0. 000)
$\Delta\ln\left(\frac{\overline{w}_{H}}{\overline{w}_{U}}\right)$ 0. 081 –0. 017 0. 574 –0. 590 0. 454 –0. 356 –0. 172 0. 058 0. 000 0. 000
(0. 022) (0. 030) (0. 028) (0. 024) (0. 016) (0. 011) (0. 002) (0. 000) (0. 000)
@@
Notes: The first column reports the observed changes in the logarithms of relative wages. The second column reports the sum of the factor complementarity effect (FCE) in the third column and the relative quantity effect (RQE) in the fourth column, or equivalently, the sum of the individual contributions of factor inputs in the fifth to last columns. The figures in the third to last columns are computed using equation (18) of Proposition 2. Standard errors in parentheses are computed using bootstrap with 500 replications.
@@
The third and fourth columns report the changes attributable to the factor complementarity effect and the relative quantity effect. The fifth to last columns report the changes attributable to each factor input. This period shows a substantial increase in ICCT capital and female skilled labor, a moderate increase in non-ICCT capital and male skilled labor, a small increase in female unskilled labor, and a small decrease in male unskilled labor. The increase in ICCT capital contributes to narrowing the gender wage gap and widening the skill wage gap. Meanwhile, the increase in male and female skilled labor contributes to widening the gender wage gap and narrowing the skill wage gap. Consequently, whether the gender (skill) wage gap narrows (widens) depends on the outcome of the race between progress in ICCT and advances in female employment (educational attainment). These results hold not only for the four wage gaps (skilled gender wage gap, unskilled gender wage gap, male skill wage gap, and female skill wage gap) reported in the first four rows, but also for the two aggregate wage gaps (aggregate gender wage gap and aggregate skill wage gap) reported in the last two rows.
The factor complementarity effect associated with a rise in ICCT capital is greater for the skilled gender wage gap and the female skill wage gap, reflecting the result that ICCT capital is significantly more complementary to female skilled labor than to the other three types of labor. If all else were held constant, the expansion of ICCT equipment during the period would have reduced the skilled gender wage gap by 88 percent and raised the female skill wage gap by 104 percent. At the same time, the relative quantity effect is greater for the skilled gender wage gap and the female skill wage gap, reflecting the fact that the rate of increase in female skilled labor is significantly greater than that in the other three types of labor. If all else were held constant, a fall in the relative quantity of male skilled labor to female skilled labor during the period would have raised the skilled gender wage gap by 95 ($(1.423-0.469)\times 100$) percent, and a rise in the relative quantity of female skilled labor to female unskilled labor would have reduced the female skill wage gap by 126 ($(\text{\textendash}1.340+0.082)\times 100$) percent. Consequently, the skilled gender wage gap did not narrow as the two large effects canceled out. Similarly, the female skill wage gap did not widen as the two large effects canceled out. Meanwhile, the unskilled gender wage gap narrowed as the factor complementarity effect exceeded the relative quantity effect. Similarly, the male skill wage gap widened as the factor complementarity effect exceeded the relative quantity effect. A rise in female skilled labor also increased the male skill wage gap as female skilled labor was more complementary to male skilled labor than to male unskilled labor. Given the results that the factor complementarity effects on the skilled gender wage gap and the female skill wage gap are so large, if there were no change in the relative supply of female skilled labor, the skilled gender wage gap would have narrowed more than the unskilled gender wage gap, and the female skill wage gap would have widened more than the male skill wage gap. Taken together, a smaller rate of decline in the gender wage gap for skilled workers than for unskilled workers and a smaller rate of increase in the skill wage gap for female workers than for male workers are attributable not to a lack of demand for female skilled labor but to an abundance in the supply of female skilled labor.
@@
Changes in the labor share can be decomposed into those in the income shares of the four types of labor $\ell_{f}$ for $f\in L=\left\{mh,fh,mu,fu\right\}$, as shown in equation (17).
$\Delta\ln\left(\frac{\sum_{f\in L}w_{f}\ell_{f}}{py}\right)=\sum_{f\in L}\left(\frac{w_{f}\ell_{f}}{\sum_{f\in L}w_{f}\ell_{f}}\right)\Delta\ln\left(\frac{w_{f}\ell_{f}}{py}\right),$ (34)
where $py=\sum_{f\in L}w_{f}\ell_{f}+r_{i}k_{i}+r_{o}k_{o}$. We evaluate the weights on the rates of changes in the income shares of each type of labor at the sample means across all countries and years. Changes in the labor share between the years 1980 and 2005 can be decomposed as
$\Delta\ln\left(\frac{\sum_{f\in L}w_{f}\ell_{f}}{py}\right)$ $=$ $\underset{\textrm{male skilled}}{\underbrace{0.165\times 0.457}}+\underset{\textrm{female skilled}}{\underbrace{0.066\times 1.230}}+\underset{\textrm{male unskilled}}{\underbrace{0.513\times\left(-0.389\right)}}+\underset{\textrm{female unskilled}}{\underbrace{0.256\times\left(-0.122\right)}}$
$=$ $0.075+0.081-0.200-0.031.$
@@
Data Total FCE RQE $\Delta\ln k_{i}$ $\Delta\ln\ell_{fh}$ $\Delta\ln\ell_{mh}$ $\Delta\ln\ell_{fu}$ $\Delta\ln\ell_{mu}$ $\Delta\ln k_{o}$
$\Delta\ln\left(\frac{w_{mh}\ell_{mh}}{py}\right)$ 0. 457 0. 438 0. 147 0. 291 0. 092 0. 050 0. 312 –0. 019 0. 003 0. 000
(0. 015) (0. 020) (0. 011) (0. 015) (0. 010) (0. 016) (0. 001) (0. 000) (0. 000)
$\Delta\ln\left(\frac{w_{fh}\ell_{fh}}{py}\right)$ 1. 230 1. 153 1. 030 0. 123 0. 976 0. 134 0. 059 –0. 018 0. 003 0. 000
(0. 047) (0. 067) (0. 051) (0. 071) (0. 060) (0. 012) (0. 002) (0. 000) (0. 000)
$\Delta\ln\left(\frac{w_{mu}\ell_{mu}}{py}\right)$ –0. 389 –0. 343 –0. 259 –0. 084 –0. 130 –0. 079 –0. 091 –0. 041 –0. 003 0. 000
(0. 011) (0. 008) (0. 009) (0. 004) (0. 003) (0. 003) (0. 001) (0. 000) (0. 000)
$\Delta\ln\left(\frac{w_{fu}\ell_{fu}}{py}\right)$ –0. 122 –0. 029 –0. 086 0. 058 –0. 060 –0. 034 –0. 042 0. 104 0. 003 0. 000
(0. 009) (0. 009) (0. 007) (0. 005) (0. 003) (0. 003) (0. 003) (0. 000) (0. 000)
$\Delta\ln\left(\frac{\sum_{f}w_{f}\ell_{f}}{py}\right)$ –0. 075 –0. 035 –0. 063 0. 028 –0. 003 –0. 032 –0. 002 0. 001 0. 000 0. 000
(0. 005) (0. 003) (0. 002) (0. 007) (0. 002) (0. 001) (0. 000) (0. 000) (0. 000)
@@
Notes: The first column reports the observed changes in the logarithms of labor shares. The second column reports the sum of the factor complementarity effect (FCE) in the third column and the relative quantity effect (RQE) in the fourth column, or equivalently, the sum of the individual contributions of factor inputs in the fifth to last columns. The figures in the third to last columns are computed using equation (15) of Proposition 1. Standard errors in parentheses are computed using bootstrap with 500 replications.
@@
Table 6 presents the quantitative contributions of specific factor inputs to changes in the income shares of the four types of labor. The first column reports the observed changes in the income shares of the four types of labor, while the second column reports the sum of the changes attributable to factor inputs. These results are obtained without adjusting any parameters to match labor shares in the aggregate economy. The observed changes in the income shares of the four types of labor align with the total contribution of factor inputs. The total contribution to changes in the income shares of male and female skilled labor are positive and statistically significant, while that in the income shares of male and female unskilled labor are negative and statistically significant.
@@
The third and fourth columns report the changes attributable to the factor complementarity effect and the relative quantity effect. The two effects work in the same direction for the income shares of male and female skilled labor and male unskilled labor. The former is much greater than the latter for the income shares of female skilled labor and male unskilled labor, although the converse is true for the income share of male skilled labor. The two effects work in opposite directions for the income share of female unskilled labor. However, the direction of the factor complementarity effect aligns with that of the actual change for the income shares of all types of labor. Overall, the factor complementarity effect is essential for understanding changes in the income shares of different types of labor. Both the factor complementarity effect and the relative quantity effect are obscured when aggregated across different types of labor, because the two effects tend to be positive for skilled labor but negative for unskilled labor.